Solve for x: 9(^2x-1)= (81^(x+1)(27^-x) I think the answer is -2??????
Answer by cngriffith(27)

9(^2x-1)= (81^(x+1)(27^-x)
(3^2)^2x-1=(3^4)^x+1(3^3)^-x Change 9,81and 27 to bases of 3
3^4x-2=3^4x+4(3^-3x) power to power: multiply exponents
3^4x-2=3^x+4 multiply like bases:add exponents
4x-2=x+4 equate exponents
3x=6
x=2